OpenStudy (anonymous):

the degree measures of the angles of triangle ABC are represented by x , 3x , and 5x-54 Find the value of x

OpenStudy (nincompoop):

how would you solve this?

OpenStudy (jdoe0001):

\(\bf x^o+(3x)^o+(5x-54)^o = \color{blue}{180}^o\quad ? \)

OpenStudy (nincompoop):

x+3x+5x-54 = 180 add 54 on both sides of the equation x+3x+5x-54+54 = 180 + 54 by adding all the x's our equation becomes 9x = 234 divide both sides by 9... do the rest of calculation